The Texas soccer league has set a budget of $1,947 for soccer balls. High-quality game balls cost $43 each while lower-quality practice balls cost $14 each. If 79 balls are to be purchased, how many balls of each type can be purchased to exactly use up the budgeted amount?
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52781)

.

Was just solved under this link

https://www.algebra.com/algebra/homework/coordinate/word/Linear_Equations_And_Systems_Word_Problems.faq.question.1192231.html

https://www.algebra.com/algebra/homework/coordinate/word/Linear_Equations_And_Systems_Word_Problems.faq.question.1192231.html



Answer by Boreal(15235)

H=x
L=79-x
43x+14(79-x)=1947
43x+1106-14x=1947
29x=841
x=29 high quality balls @$43=$1247
79-x=50 lower quality balls @$14=$700